OpenStudy (anonymous):

Evaluate 12(w - 5z), given values w = 16, z = 3. 12 121 396 -396

OpenStudy (experimentx):

12

OpenStudy (bahrom7893):

12(16-15)=12*1=12

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

12(w - 5z)= 12(16 - 5(3)) = 12(16-15) = 12(1) =12

OpenStudy (anonymous):

12